OpenStudy (anonymous):

A neighborhood has provided trash cans to residents to dispose of their waste. All trash cans have a diameter of 1 meter and a height of 0.75 meters. What is the minimum number of trash cans required to dispose of 10 cubic meters of waste? 16 17 50 51

OpenStudy (ghazi):

hey...volume of one trash can= 3.14*(1/2)^2*0.75=0.588 cubic meter therefore for one cubic meter you need 1/0.588 can ans for 10 cubic meter you need = 10*(1/0.588)=17 cans

OpenStudy (ghazi):

@devibabyy66 clear?

OpenStudy (anonymous):

yeah thanks @ghazi :)

OpenStudy (ghazi):

:) YW